3x+6 over 4x+12 divided by x squared minus 4 over x-3
Answer by jorel1380(3719)

((3x+6)/(4x+12))/((x2-4)/(x-3))=3(x+2)/4(x+3)x(x-3)/(x+2)(x-2)=
3/4(x+3)x(x-3)/(x-2)=3(x-3)/4(x+3)(x-2)=3x-9/4x2+4x-24..